# Scalable Simulation of Fermionic Encoding Performance on Noisy Quantum Computers

### v1

@@ I Introduction
It is widely believed that chemical and materials simulations will be a high-impact application of future quantum computers [1, 2, 3, 4, 5]. For these applications the relevant degrees of freedom are often fermionic and therefore it is necessary to represent fermions using the building blocks of a quantum computer: qubits. The first such representation (or encoding) was the Jordan-Wigner (JW) encoding [6], which has the advantage of efficiency in that it uses $n$ qubits to represent $n$ fermions. However, in the JW encoding local fermionic operators are often represented by non-local and high-weight qubit operators, which can require many gates to implement on hardware.
A variety of other fermionic encodings have been proposed in more recent times [7, 8, 9, 10, 11, 12, 13], some of which map local fermionic operators to local qubit operators at the cost of requiring more than one qubit per encoded fermion. Some of these encodings have the property that the encoded states are contained in a subspace of the full Hilbert space, which allows us to detect errors by checking whether a state resides in this subspace. And in certain encodings such as the Derby-Klassen (DK) encoding [14], the subspace is stabilized by local Pauli operators similar to those of topological codes [15, 16] and we can therefore detect errors by measuring these local operators.
Small-scale demonstrations of quantum chemistry have been carried out on quantum computers with promising results [17, 18, 19, 20]. But achieving quantum advantage in these tasks likely requires a larger number of qubits and deeper circuits than are possible with today’s devices [21]. Furthermore, predicting the performance of quantum computers in these tasks is fraught with difficulty due to our inability of simulating large and deep quantum circuits. However certain sub-theories of quantum mechanics, such as stabilizer quantum mechanics [22], are classically simulable but even so exhibit quantum phenomena such as entanglement and superposition. Indeed, stabilizer simulations are widely used to benchmark the performance of quantum error-correcting (QEC) codes under Pauli error models, and these simulations are believed to give a reliable indication of performance.
@@ II.4 Derby-Klassen encoding
The DK encoding has an encoded Majorana fermion for every vertex of $\mathcal{L}_{v}$ and is customarily defined in terms of a convenient Majorana operator basis:
@@ III.4 Readout
Since, by definition, fermionic encodings preserve commutation relations between the terms of the Hamiltonian, there must be the same number of sets of commuting subsets of terms in the JW and ternary tree encodings. However, there is no guarantee that those terms will be qubit-wise commuting, and we therefore may require a diagonalization circuit to allow for simultaneous measurement. In a square-lattice Fermi-Hubbard Hamiltonian encoded with the JW encoding, it is known how to extract the energy with 5 measurements using the procedure outlined by [47]. This procedure employs a simple, but non-Clifford, diagonalization circuit for the two parts of the hopping term. To the best of our knowledge, no efficient measurement procedure is known for ternary tree encoding.
@@ V Conclusion
In this work we used modern stabilizer simulation tools to estimate the performance of fermionic encodings on near-term devices. We were able to assess the performance of encodings in regimes that are potentially impractical for classical computation [65], reaching lattice sizes of up to $18\times 18$ and up to 40 Trotter steps. To achieve this our error model was necessarily restricted to stochastic Pauli noise, but realistic noise in devices can be tailored to this form at little or no experimental cost.
@@
In our simulations, we observed that the asymptotic efficiency of the TT encoding cannot compete with depth reductions provided with specialized methods such the JW encoding with fermionic swap netwoks and the DK encoding with error mitigation. We also observed that although the DK encoding with error detection provides an accuracy improvement, the returns are diminishing due to the high sampling requirements. The most promising regime for the DK encoding is the aspirational regime ($p=0.01\%$). However, even in this case the DK encoding exceeds our sample budget of 100,000 shots at 10 Trotter steps for the $12\times 12$ lattice or 6 Trotter steps for the $16\times 16$ lattice. Other techniques we tested such as flag qubits and virtual quantum error detection (VQED) did not show improvements in accuracy or sampling requirements. Furthermore, in the aspirational regime, the JW encoding with fermionic swap networks and global parity postselection also provides relatively low-error results.

### v2

@@ I Introduction
It is widely believed that chemical and materials simulations will be a high-impact application of future quantum computers [1, 2, 3, 4, 5]. For these applications the relevant degrees of freedom are often fermionic and therefore it is necessary to represent fermions using the building blocks of a quantum computer: qubits. The first such representation (or encoding) was the Jordan-Wigner (JW) encoding [6], which has the advantage of efficiency in that it uses $n$ qubits to represent $n$ fermionic modes. However, in the JW encoding local fermionic operators are often represented by non-local and high-weight qubit operators, which can require many gates to implement on hardware.
A variety of other fermionic encodings have been proposed in more recent times [7, 8, 9, 10, 11, 12, 13, 14, 15], some of which map local fermionic operators to local qubit operators at the cost of requiring more than one qubit per encoded fermion. Some of these encodings have the property that the encoded states are contained in a subspace of the full Hilbert space, which allows us to detect errors by checking whether a state resides in this subspace. And in certain encodings such as the Derby-Klassen (DK) encoding [16], the subspace is stabilized by local Pauli operators similar to those of topological codes [17, 18] and we can therefore detect errors by measuring these local operators.
Small-scale demonstrations of quantum chemistry have been carried out on quantum computers with promising results [19, 20, 21, 22]. But achieving quantum advantage in these tasks likely requires a larger number of qubits and deeper circuits than are possible with today’s devices [23]. Furthermore, predicting the performance of quantum computers in these tasks is fraught with difficulty due to our inability of simulating large and deep quantum circuits. However, certain sub-theories of quantum mechanics, such as stabilizer quantum mechanics [24], are classically simulable but even so exhibit quantum phenomena such as entanglement and superposition. Indeed, stabilizer simulations are widely used to benchmark the performance of quantum error-correcting (QEC) codes under Pauli error models, and these simulations are believed to give a reliable indication of performance.
@@ II.4 Derby-Klassen encoding
The DK encoding has an encoded fermionic mode for every vertex of $\mathcal{L}_{v}$ and is customarily defined in terms of a convenient Majorana operator basis:
@@ III.4 Readout
In a square-lattice Fermi-Hubbard Hamiltonian encoded with the JW encoding, it is known how to extract the energy with 5 measurements using the procedure outlined by [49]. This procedure employs a simple, but non-Clifford, diagonalization circuit for the two parts of the hopping term. To the best of our knowledge, no efficient measurement procedure is known for ternary tree encoding.
@@ V Conclusion
In this work we used modern stabilizer simulation tools to estimate the performance of fermionic encodings on near-term devices. We were able to assess the performance of encodings in regimes that are potentially impractical for classical computation [67, 68], reaching lattice sizes of up to $18\times 18$ and up to 40 Trotter steps. To achieve this our error model was necessarily restricted to stochastic Pauli noise, but realistic noise in devices can be tailored to this form at little or no experimental cost.
@@
In our simulations, we observed that the optimal average operator weight of the TT encoding cannot compete with depth reductions provided with specialized methods such as the JW encoding with fermionic swap netwoks and the DK encoding with error mitigation. We also observed that although the DK encoding with error detection provides an accuracy improvement, the returns are diminishing due to the high sampling requirements. The most promising regime for the DK encoding is the aspirational regime ($p=0.01\%$). However, even in this case the DK encoding exceeds our sample budget of 100,000 shots at 10 Trotter steps for the $12\times 12$ lattice or 6 Trotter steps for the $16\times 16$ lattice. Other techniques we tested such as flag qubits and virtual quantum error detection (VQED) did not show improvements in accuracy or sampling requirements. Furthermore, in the aspirational regime, the JW encoding with fermionic swap networks and global parity postselection also provides relatively low-error results.